For the line 3x-7y=21 state the x-intercept
For the line 3x-7y=21 state the y-intercept
For the line 3x-7y=21 state the slope
For the line 3x-7y=21 draw the graph

Answer by ewatrrr(24785)

 
Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
For the line 3x-7y=21 state the x-intercept )when y = 0, x = 7 Pt(7,0)
3x-7y=21 OR y = (3/7)x - 3
For the line 3x-7y=21 state the y-intercept Pt(0,-3)
For the line 3x-7y=21 state the slope slope m = (3/7)